A bag of cement weighing 325 N hangs in equilibrium from three wires. Two of the wires make make angles u1 60.0 degrees and u2 40.0 degrees with the horizontal. Assuming the system is in equilibrium; find the tensions T 1 , T 2 , and T 3 in the wires
1
Expert's answer
2020-09-22T15:36:12-0400

"T_3=325(N)"


"T_1\cos60°=T_2\cos40°\to T_2=T_1\cos60°/\cos40°"


"T_1\sin60°+T_2\sin40°=T_3"


"T_1\sin60°+T_1\cos60°\sin40°/\cos40°=T_3"


"T_1=\frac{T_3}{\sin60°+\cos60°\sin40°/\cos40°}="


"=\frac{325}{\sin60°+\cos60°\sin40°/\cos40°}=253(N)"


"T_2=T_1\cos60°/\cos40°="


"=253\cdot\cos60°/\cos40°=165(N)"